OpenStudy (anonymous):

domain ln(x^2 −10x) ?

OpenStudy (anonymous):

if you have ln(n), what must n be?

OpenStudy (anonymous):

base?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

ok so f(x) is ln(x^2 −10x) and i found f'(x)=2(x-5)/x(x-10)

OpenStudy (anonymous):

now the questions asks for domain for f

OpenStudy (anonymous):

of*

OpenStudy (anonymous):

you don't use derivative to find the domain of f

OpenStudy (anonymous):

can you take ln(-4), ln(-3) , ln(-0.2), ...?

OpenStudy (anonymous):

that was the first part of question

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how about ln(0) ?

OpenStudy (anonymous):

so have you found the domain of f yet?

OpenStudy (anonymous):

yupp i got it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(-\[(-\infty, 0) and (10, \infty)\]

OpenStudy (anonymous):

but what is the correct way to write it?